# oMEGACat II - Photometry and proper motions for 1.4 million stars in Omega Centauri and its rotation in the plane of the sky

### v1

@@ Abstract
Omega Centauri ($\omega$ Cen) is the most massive globular cluster of the Milky Way. It is thought to be the nucleus of an accreted dwarf galaxy because of its high mass and its complex stellar populations. To decipher its formation history and study its dynamics, we created the most comprehensive kinematic catalog for its inner region, by analyzing both archival and new Hubble Space Telescope (HST) data. Our catalog contains 1 395 781 proper-motion measurements out to the half-light radius of the cluster ($\sim 5.0\arcmin$) and down to $m_{\rm F625W}\approx 25$. The typical baseline for our proper-motion measurements is 20 years, leading to a median 1D proper motion precision of $\sim$11 µas yr-1 for stars with $m_{\rm F625W}\approx 18$ mag, with even better precision ($\sim$6.6 µas yr-1) achieved in the extensively observed centermost ($r<1.5\arcmin$) region. In addition to our astrometric measurements, we also obtained precise HST photometry in seven filters spanning from the ultraviolet to the near-infrared. This allows detailed color-magnitude-diagram studies and to separate the multiple stellar populations of the cluster. In this work, we describe the data reduction used to obtain both the photometric and the proper-motion measurements. We also illustrate the creation and the content of our catalog, which is made publicly available. Finally, we present measurements of the plane-of-sky rotation of $\omega$ Cen in the previously unprobed inner few arcminutes and a precise measurement of the inclination $i=(43.6\pm 1.5)^{\circ}$.
@@ 1.1 The accretion history of the Milky Way
The largest of the recently discovered mergers is the Gaia Enceladus event, a merger $\sim$10 Gyrs ago with a satellite galaxy with a stellar mass of $6\times 10^{8}$ M (Belokurov et al., 2018; Haywood et al., 2018), similar to the present-day mass of the Small Magellanic Cloud. There are also signs of other smaller accretion events such as Sequoia (Myeong et al., 2019) and the Helmi-Streams (Helmi et al., 1999). An example, where accretion is still ongoing, is the Sagittarius Dwarf Galaxy (Ibata et al., 1997; Laporte et al., 2018). The central regions of Sagittarius remain bound, including the nuclear star cluster of the galaxy, M54 (Alfaro-Cuello et al., 2019, 2020; Kacharov et al., 2022).
@@
Besides these peculiar stellar populations, there is also kinematic evidence supporting the stripped nucleus scenario: van de Ven et al. (2006) found evidence for the presence of a central stellar disk and a preference for tangential orbits in the outer parts. More recently both kinematic and chemical associations with stellar streams such as the Fimbulthul stream have been found in e.g. Majewski et al. (2012), Ibata et al. (2019), and Limberg et al. (2022). Both the Sequoia and the Gaia-Enceladus/Sausage progenitors have been discussed as potential former host-galaxies of $\omega$ Cen (Myeong et al., 2019; Massari et al., 2019; Forbes, 2020; Pfeffer et al., 2021).
@@ 2.1 Other proper motion catalogs
In comparison, the Hubble Space Telescope proved to be the perfect tool for crowded field astrometry. Its high resolution and well-characterized, stable point spread function allow individual astrometric measurements with 0.4 mas precision (Anderson & King, 2006; Bellini et al., 2011). Thanks to its high sensitivity, very faint stars can also be studied. For $\omega$ Cen, the main limitation is the comparatively small field of view of its main imaging instruments ($3\farcm 3\times 3\farcm 3$ for ACS/WFC, $2\farcm 7\times 2\farcm 7$ for WFC3/UVIS); the existing HST proper-motion catalogs cover an area of only one or two HST pointings. The first HST proper motion study of $\omega$ Cen’s innermost region was done in Anderson & van der Marel (2010), with an additional field South-East of the center. The measurements in this area were significantly improved in Bellini et al. (2014) and published along an extensive photometric catalog in Bellini et al. (2017a). This most recent public catalog covers the core region out to a radius of $\farcm$7 and has a maximum temporal baseline of 12 years. One other notable work based on HST observations of the center of $\omega$ Cen was the detection of astrometric acceleration by dark companions presented in Platais et al. (2023), however the astrometric catalog has not been made public. Other HST fields at larger radii have been analyzed in Bellini et al. (2018b) ($r\sim 17\arcmin\sim 3.5r_{\rm HL}$) and Scalco et al. (2021) ($r\sim 12\arcmin\sim 2.5r_{\rm HL}$).
@@ 2.2 Other HST photometry catalogs of $\omega$ Cen
In the past years, several photometric catalogs based on HST imaging have been published for $\omega$ Cen with various science goals. Some of them were published together with the astrometric catalogs already mentioned in the previous section. The first HST based photometric catalog of $\omega$ Cen was created as part of the the “The ACS Survey of Globular Clusters” and is described in Anderson et al. (2008). A much larger catalog was then published by Anderson & van der Marel (2010), covering a grid of $3\times 3$ ACS/WFC pointings, giving an on-sky extent of $10\arcmin\times 10\arcmin$ and containing deep photometry for 2×$\times 10^{6}$ stars in the F435W and F625W filters. The data used for this study - observed in 2002 - marks the first epoch for most of our proper-motion measurements. With the installation of the WFC3/UVIS instrument, a new range of UV filters became available. They have been used to study multiple populations, for example in Bellini et al. (2010) and Bellini et al. (2013). The deepest photometry for $\omega$ Cen has been obtained for a Large Project (Milone et al., 2017b) with the goal of studying the stars at the faint end of the main sequence. Several astrometric and photometric catalogs based on this data have been published (Bellini et al., 2018b; Libralato et al., 2018a; Scalco et al., 2021; Gerasimov et al., 2022). The most recent catalog for the core of $\omega$ Cen was published by Bellini et al. (2017a), containing the most comprehensive set of filters (18 WFC3/UVIS filters and 8 WFC3/IR filters) and the same state-of-the-art photometry software as in this work. This catalog is limited to the centermost region of $\omega$ Cen with $r\leq 2\farcm 5$.
@@ 3 Dataset
For our study, we used imaging data obtained with the Advanced Camera for Surveys (ACS) Wide Field Channel and the Wide Field Camera 3 UVIS Channel. The data from both of these instruments are similar: both instruments have a mosaic of two 2048$\times$4096 pixel CCD detectors with a narrow chip gap, giving approximately a square footprint on the sky. The ACS/WFC, installed during Service Mission 3B, has a nominal pixel scale of 50 mas pixel-1, giving a field of view of $3\farcm 3\times 3\farcm 3$. WFC3/UVIS was installed during Service Mission 4 and has a slightly higher resolution with a pixel scale of 40 mas pixel-1, resulting in a field of view of $2\farcm 7\times 2\farcm 7$.
@@
In total, we reduced 323 images taken with the Advanced Camera for Surveys (ACS) Wide Field Channel, and $1$ images taken with the Wide Field Camera 3 UVIS Channel, including both archival data and data from a new, dedicated program (GO-16777, PI: A. Seth). However, not all filters are suitable for high-precision astrometry, due to the unavailability of dedicated high-precision geometric distortion corrections. Proper motions are therefore based on a subset of the data including 196 ACS/WFC and 476 WFC3/UVIS exposures. For the photometric catalogs we restricted ourselves to the six filters with the widest field coverage (WFC3/UVIS F275W, F336W, F814W; ACS/WFC F435W, F625W, F658N). These 6 filters fill the half-light radius with only minimal gaps. In addition, we also included the WFC3/UVIS F606W filter which has only been used in the central region of $\omega$ Cen. Due to the large number of calibration observations ($N_{used}=184$) in this filter it provides excellent photometric quality out to $r\sim 2.5\arcmin$. Footprints of the utilized observations can be found in Figure 1. Table 2 lists all filters and the number of images used for the creation of our catalogs. Although available in the archive, we did not make use of any WFC3/IR images, as they are less useful for high-precision astrometry due to their relatively large pixel size (130 mas pixel-1). A state-of-the-art reduction of the WFC3/IR data can be found in Bellini et al. (2017a). A detailed list of all program IDs, filters, and exposure times used for our analysis is shown in Tables A and A in the appendix. In addition, all the HST data used in this paper can be found under the following DOI in the Mikulski Archive for Space Telescopes (MAST): http://dx.doi.org/10.17909/26qj-g090 (catalog 10.17909/26qj-g090).
@@
Table 2: List of all filters used for the creation of our astro-photometric catalog. Filters marked with have only been used for photometry, but not for the astrometric measurements.
Instrument Filter Nexposures
ACS/WFC F435W 69
ACS/WFC F475W 7
ACS/WFC F [^5] W 4
ACS/WFC F606W 35
ACS/WFC F625W 40
ACS/WFC F658N 39
ACS/WFC F775W 8
ACS/WFC F814W 33
WFC3/UVIS F275W 85
WFC3/UVIS F336W 106
WFC3/UVIS F390W 15
WFC3/UVIS F438W 49
WFC3/UVIS F [^5] W 25
WFC3/UVIS F606W 184
WFC3/UVIS F775W 18
WFC3/UVIS F814W 79
@@ 4.1 First photometry iteration with hst1pass
hst1pass uses the library effective PSF (ePSF) models described in Anderson & King (2006). When available, we used the state-of-the-art focus-diverse ePSF models described in Bellini et al. (2018a) for ACS/WFC and Anderson (2018) for WFC3/UVIS. hst1pass further improves the library ePSF to better match those of each individual image. This is necessary due to the time variability of the HST PSF caused by telescope breathing and focus variations. We saved these perturbed ePSF models for each image, as we need them later for the second photometry stage.
@@ 5.2 Iterative Proper Motion Determination
Proper motions are measured using the method developed in Bellini et al. (2014) and improved in Bellini et al. (2018b); Libralato et al. (2018b), and we refer to those publications for detailed descriptions of the procedure. In short, the program treats each individual image catalog as a stand-alone epoch. As a first step, all these geometric-distortion-corrected [^5] Our software uses the library geometric distortion corrections from Anderson (2003) (ACS), Bellini & Bedin (2009), and Bellini et al. (2011) (UVIS), with an additional look-up table for ACS observations post HST-SM4 (2009). We noticed that the ACS/WFC distortion has worsened over time. For this reason, we made additional, time-depend table-of-residuals corrections for the latest observations (epoch $>$ 2018) with the ACS/WFC detector following the prescriptions in Bellini et al. (2011). individual-image catalogs are transformed to the reference frame using linear six-parameter transformations. These transformations are determined individually for the 4 amplifiers used to read out the images (i.e. the single image catalogs are split into 4 quadrants, corresponding to the 4 amplifiers reading out the detectors of the instruments) to mitigate potential amplifier-based systematic effects. In the final iteration, the transformations are determined individually for each star using its 100 closest reference stars. The transformed positions of each star are fitted with a straight line in both the $x$ and $y$ directions, to directly fit the two proper-motion components. The fit takes into account the magnitude-dependent astrometric errors and has several stages of outlier rejection. Proper motions are iteratively improved by removing non-members from the set of reference stars based on their position in the color-magnitude diagram and in later iterations also the vector point diagram (Figure 7), or because of spurious proper-motion measurements. In another iterative loop, the crossmatch and the transformations to the master frame positions are improved by using the proper motions to propagate the reference stars to the same epoch as the individual image catalogs. As all reference stars are bona-fide cluster members, our proper motions are determined relative to the bulk motion of $\omega$ Cen. From 1 482 835 stars that have been measured at least twice, 1 395 781 pass the iteration process and have a high-precision proper-motion measurement. The vast majority of our proper motions (1 102 818) have a temporal baseline longer than 20 years. The median number of individual astrometric measurements used for the proper-motion determination is 17, it is lowest in the outskirts of our field of view and quickly increases towards the center. In the very center and in the best-covered magnitude range ($m_{\rm F625W}=17.5$ to $m_{\rm F625W}=22.0$), many stars have more than 400 individual measurements (with a maximum of 467), leading to a median proper-motion error of only 6.6 µas yr-1 ($\sim$0.15 km s-1 at the distance of $\omega$ Cen) with individual stars reaching as low as 3.3 µas yr-1. The field dependence of the temporal baseline and the number of available measurements and their effect on the proper-motion error are presented in Figures 4 and 5.

### v2

@@ Abstract
Omega Centauri ($\omega$ Cen) is the most massive globular cluster of the Milky Way. It is thought to be the nucleus of an accreted dwarf galaxy because of its high mass and its complex stellar populations. To decipher its formation history and study its dynamics, we created the most comprehensive kinematic catalog for its inner region, by analyzing both archival and new Hubble Space Telescope (HST) data. Our catalog contains 1 395 781 proper-motion measurements out to the half-light radius of the cluster ($\sim 5.0\arcmin$) and down to $m_{\rm F625W}\approx 25$ mag. The typical baseline for our proper-motion measurements is 20 years, leading to a median 1D proper motion precision of $\sim$11 µas yr-1 for stars with $m_{\rm F625W}\approx 18$ mag, with even better precision ($\sim$6.6 µas yr-1) achieved in the extensively observed centermost ($r<1.5\arcmin$) region. In addition to our astrometric measurements, we also obtained precise HST photometry in seven filters spanning from the ultraviolet to the near-infrared. This allows detailed color-magnitude-diagram studies and to separate the multiple stellar populations of the cluster. In this work, we describe the data reduction used to obtain both the photometric and the proper-motion measurements. We also illustrate the creation and the content of our catalog, which is made publicly available. Finally, we present measurements of the plane-of-sky rotation of $\omega$ Cen in the previously unprobed inner few arcminutes and a precise measurement of the inclination $i=(43.9\pm 1.3)^{\circ}$.
@@ 1.1 The accretion history of the Milky Way
The largest of the recently discovered mergers is the Gaia Enceladus event, a merger $\sim$10 Gyrs ago with a satellite galaxy with a stellar mass of $6\times 10^{8}$ M (Belokurov et al., 2018; Haywood et al., 2018), similar to the present-day mass of the Small Magellanic Cloud. There are also signs of other smaller accretion events such as Sequoia (Myeong et al., 2019),the Helmi-Streams (Helmi et al., 1999) or the Pontus merger Malhan et al. (2022); Malhan (2022). An example, where accretion is still ongoing, is the Sagittarius Dwarf Galaxy (Ibata et al., 1997; Laporte et al., 2018). The central regions of Sagittarius remain bound, including the nuclear star cluster of the galaxy, M54 (Alfaro-Cuello et al., 2019, 2020; Kacharov et al., 2022).
@@
Besides these peculiar stellar populations, there is also kinematic evidence supporting the stripped nucleus scenario: van de Ven et al. (2006) found evidence for the presence of a central stellar disk and a preference for tangential orbits in the outer parts. More recently, both kinematic and chemical associations with stellar streams such as the Fimbulthul stream have been found in e.g. Majewski et al. (2012), Ibata et al. (2019), and Limberg et al. (2022). Another approach is taken in Marks et al. (2022), in which a connection between the low retrograde binary fraction in the Milky Way and the star formation conditions in $\omega$ Cen’s progenitor is studied. Both the Sequoia and the Gaia-Enceladus/Sausage progenitors have been discussed as potential former host-galaxies of $\omega$ Cen (Myeong et al., 2019; Massari et al., 2019; Forbes, 2020; Pfeffer et al., 2021).
@@ 2.1 Other proper motion catalogs
In comparison, the Hubble Space Telescope proved to be the perfect tool for crowded field astrometry. Its high resolution and well-characterized, stable point spread function allow individual astrometric measurements with 0.4 mas precision (Anderson & King, 2006; Bellini et al., 2011). Thanks to its high sensitivity, very faint stars can also be studied. For $\omega$ Cen, the main limitation is the comparatively small field of view of its main imaging instruments ($3\farcm 3\times 3\farcm 3$ for ACS/WFC, $2\farcm 7\times 2\farcm 7$ for WFC3/UVIS); the existing HST proper-motion catalogs cover an area of only one or two HST pointings. The first HST proper motion study of $\omega$ Cen’s innermost region was done in Anderson & van der Marel (2010), with an additional field South-East of the center. The measurements in this area were significantly improved in Bellini et al. (2014) and published along an extensive photometric catalog in Bellini et al. (2017a). This most recent public catalog covers the core region out to a radius of $\farcm$7 and has a maximum temporal baseline of 12 years. One other notable work based on HST observations of the center of $\omega$ Cen was the detection of astrometric acceleration by dark companions presented in Platais et al. (2023), however the astrometric catalog has not been made public. Other HST fields at larger radii have been analyzed in Bellini et al. (2018b) ($r\sim 17\arcmin\sim 3.5r_{\rm HL}$) and Scalco et al. (2021) ($r\sim 12\arcmin\sim 2.5r_{\rm HL}$). Due to the lower stellar density at these radii and the long exposure times, those fields mark the deepest observations of $\omega$ Cen at the time of writing, reaching magnitudes of $m_{\rm F606W}\sim 27$.
@@ 2.2 Other HST photometry catalogs of $\omega$ Cen
In the past years, several photometric catalogs based on HST imaging have been published for $\omega$ Cen with various science goals. Some of them were published together with the astrometric catalogs already mentioned in the previous section. The first HST based photometric catalog of $\omega$ Cen was created as part of the “The ACS Survey of Globular Clusters” and is described in Anderson et al. (2008). A much larger catalog was then published by Anderson & van der Marel (2010), covering a grid of $3\times 3$ ACS/WFC pointings, giving an on-sky extent of $10\arcmin\times 10\arcmin$ and containing deep photometry for 2×$\times 10^{6}$ stars in the F435W and F625W filters. The data used for this study - observed in 2002 - marks the first epoch for most of our proper-motion measurements. With the installation of the WFC3/UVIS instrument, a new range of UV filters became available. They have been used to study multiple populations, for example in Bellini et al. (2010) and Bellini et al. (2013). The deepest photometry for $\omega$ Cen has been obtained for a Large Project (Milone et al., 2017b) with the goal of studying the stars at the faint end of the main sequence. Several astrometric and photometric catalogs based on this data have been published (Bellini et al., 2018b; Libralato et al., 2018a; Scalco et al., 2021; Gerasimov et al., 2022). The most recent catalog for the core of $\omega$ Cen was published by Bellini et al. (2017a), containing the most comprehensive set of filters (18 WFC3/UVIS filters and 8 WFC3/IR filters) and the same state-of-the-art photometry software as in this work. This catalog is limited to the centermost region of $\omega$ Cen with $r\leq 2\farcm 5$.
@@ 3 Dataset
For our study, we used imaging data obtained with the Advanced Camera for Surveys (ACS) Wide Field Channel and the Wide Field Camera 3 UVIS Channel. The data from both of these instruments are similar: both instruments have a mosaic of two 2048$\times$4096 pixel CCD detectors with a narrow chip gap, giving approximately a square footprint on the sky. The ACS/WFC, installed during Service Mission 3B, has a nominal pixel scale of around 50 mas pixel-1, giving a field of view of $3\farcm 3\times 3\farcm 3$. WFC3/UVIS was installed during Service Mission 4 and has a slightly higher resolution with a pixel scale of around 40 mas pixel-1, resulting in a field of view of $2\farcm 7\times 2\farcm 7$.
@@
In total, we reduced 236 images taken with the Advanced Camera for Surveys (ACS) Wide Field Channel, and 561 images taken with the Wide Field Camera 3 UVIS Channel, including both archival data and data from a new, dedicated program (GO-16777, PI: A. Seth). However, not all filters are suitable for high-precision astrometry, due to the unavailability of dedicated high-precision geometric distortion corrections. Proper motions are, therefore, based on a subset of the data, including 196 ACS/WFC and 476 WFC3/UVIS exposures. For the photometric catalogs, we restricted ourselves to the six filters with the widest field coverage (WFC3/UVIS F275W, F336W, F814W; ACS/WFC F435W, F625W, F658N). These 6 filters fill the half-light radius with only minimal gaps. In addition, we also included the WFC3/UVIS F606W filter which has only been used in the central region of $\omega$ Cen. Due to the large number of calibration observations ($N_{used}=184$) in this filter it provides excellent photometric quality out to $r\sim 2.5\arcmin$. Footprints of the utilized observations can be found in Figure 1. Table 2 lists all filters and the number of images used for the creation of our catalogs. We note that, while the ACS High Resolution Channel is principally suitable for high-precision astrometry, there are no usable observations within the field covered in this study. Although available in the archive, we also did not make use of any WFC3/IR images, as they are less useful for high-precision astrometry due to their relatively large pixel size (130 mas pixel-1). A state-of-the-art reduction of the WFC3/IR data can be found in Bellini et al. (2017a). A detailed list of all program IDs, filters, and exposure times used for our analysis is shown in Tables A and A in Appendix A. In addition, all the HST data used in this paper can be found under the following DOI in the Mikulski Archive for Space Telescopes (MAST): http://dx.doi.org/10.17909/26qj-g090 (catalog 10.17909/26qj-g090).
@@
Table 2: List of all filters used for the creation of our astro-photometric catalog. In the fourth column we state whether a filter has been used only for photometry (phot.), astrometry (astro.), or both.
Instrument Filter Nexp. Usage
ACS/WFC F435W 69 astro. & phot.
ACS/WFC F475W 7 astro. only
ACS/WFC F [^5] W 4 astro. only
ACS/WFC F606W 35 astro. only
ACS/WFC F625W 40 astro. & phot.
ACS/WFC F658N 39 phot. only
ACS/WFC F775W 8 astro. only
ACS/WFC F814W 33 astro. only
WFC3/UVIS F275W 85 phot. only
WFC3/UVIS F336W 106 astro. & phot.
WFC3/UVIS F390W 15 astro. only
WFC3/UVIS F438W 49 astro. only
WFC3/UVIS F [^5] W 25 astro. only
WFC3/UVIS F606W 184 astro. & phot.
WFC3/UVIS F775W 18 astro. only
WFC3/UVIS F814W 79 astro. & phot.
@@ 4.1 First photometry iteration with hst1pass
hst1pass uses the library effective PSF (ePSF) models described in Anderson & King (2006). When available, we used the state-of-the-art focus-diverse ePSF models described in Bellini et al. (2018a) for ACS/WFC and Anderson (2018) for WFC3/UVIS. hst1pass further improves the library ePSF to better match those of each individual image. This is necessary due to the time variability of the HST PSF caused by telescope breathing and focus variations. For typical observations, the central values of the PSF change by around 5% (RMS over the full field), although in some rare cases, this change can be up to 25%. We saved these perturbed ePSF models for each image, as we need them later for the second photometry stage.
@@ 5.2 Iterative Proper Motion Determination
Proper motions are measured using the method developed in Bellini et al. (2014) and improved in Bellini et al. (2018b); Libralato et al. (2018b), and we refer to those publications for detailed descriptions of the procedure. Proper motions are measured relative to a subset of well-measured cluster stars. The set of reference stars is iteratively improved. In the first iteration, it is based on photometric-quality indicators and the stars’ position in the color-magnitude diagram alone. Once proper motions become available, non-members are also removed based on their position in the vector-point diagram (Figure 7) or if they have a spurious proper motion measurement.
@@
The program treats each individual image catalog as a stand-alone epoch. As a first step, all these geometric-distortion-corrected [^5] Our software uses the library geometric distortion corrections from Anderson (2003) (ACS), Bellini & Bedin (2009), and Bellini et al. (2011) (UVIS), with an additional look-up table for ACS observations post HST-SM4 (2009). We noticed that the ACS/WFC distortion has worsened over time. For this reason, we made additional, time-depend table-of-residuals corrections for the latest observations (epoch $>$ 2018) with the ACS/WFC detector following the prescriptions in Bellini et al. (2011). individual-image catalogs are transformed to the reference frame using linear six-parameter transformations. These transformations are determined individually for the 4 amplifiers used to read out the images (i.e. the single image catalogs are split into 4 quadrants, corresponding to the 4 amplifiers reading out the detectors of the instruments) to mitigate potential amplifier-based systematic effects. In the final iteration, the transformations are determined individually for each star using its 100 closest reference stars. The transformed positions of each star are fitted with a straight line in both the $x$ and $y$ directions, to directly fit the two proper-motion components. The fit takes into account the magnitude-dependent astrometric errors and has several stages of outlier rejection.
In another iterative loop, the crossmatch and the transformations to the master frame positions are improved by using the proper motions to propagate the reference stars to the same epoch as the individual image catalogs. In total, 103 616 339 individual position measurements were used for the proper-motion measurements, making this one of the largest astrometric datasets of all times.
Some detections that appeared as two separate sources during the crossmatch could be reassigned to a single source using the proper motions. Therefore, the final number of entries in our catalog is 1 475 096, slightly lower than the 1 482 835 sources measured at least twice during the crossmatch.
In total, 1 395 781 individual sources pass the iterative process and have a high-precision proper-motion measurement.
The vast majority of our proper motions (1 102 818) have a temporal baseline longer than 20 years. The median number of individual astrometric measurements used for the proper-motion determination is 17, it is lowest in the outskirts of our field of view and quickly increases towards the center. In the very center and in the best-covered magnitude range ($m_{\rm F625W}=17.5$ to $m_{\rm F625W}=22.0$), many stars have more than 400 individual measurements (with a maximum of 467), leading to a median proper-motion error of only 6.6 µas yr-1 ($\sim$0.15 km s-1 at the distance of $\omega$ Cen) with individual stars reaching as low as 3.3 µas yr-1. The field dependence of the temporal baseline and the number of available measurements and their effect on the proper-motion error are presented in Figures 4 and 5.